OpenStudy (anonymous):

Solve By Completing the Square. Give Exact Answer Only. 2w^2 - 3w = 4

OpenStudy (anonymous):

\[2w^2-3w=4\] Divide both sides by 2 to get the coefficient of w^2 to 1. \[w^2-\frac{3}{2}w=2\] Completing the square is essentially \[(x-a)^2 = x^2-2ax+a^2\] what you have is a part of the right side of that expression. \[(x-a)^2 = x^2-2ax+a^2=x^2-\frac{3}{2}x+a^2\] So now, you have to find what that a value is. looking at the \[-2ax=-\frac{3}{2}x\], you can find that the a value is 3/4. Plug it back into the equation and you get \[(x-\frac{3}{4})^2 = x^2-\frac{3}{2}x+\frac{9}{4}\] Notice how you have an extra term in there. To make up for adding that term, simply subtract it from the expression. \[(x-\frac{3}{4})^2-\frac{9}{4}=x^2-\frac{3}{2}x+\frac{9}{4}-\frac{9}{4}=x^2-\frac{3}{2}x\] So your final equation just needs to simplify \[(x-\frac{3}{4})^2-\frac{9}{4}=2\]